An ideal step up transformer is primary coil has 500 turns and it’s secondary coil has 15,000 turns. The primary EMF is 120V. What is the EMF of the secondary?


1
Expert's answer
2022-02-07T08:43:00-0500

"\frac{\varepsilon_1}{\varepsilon_2}=\frac {N_1 }{N_2},"

"\varepsilon_2=\varepsilon_1\frac{N_2}{N_1}=3.6~kV."